Geometry: Floor plans for a building have the four corners of a room located at the points (2,3), (11,6), (-3,18), and (8,21). Determine whether the side through the points (2,3) and (11,6) is perpendicular ot the side through the points (2,3) and (-3,18).
Answer by venugopalramana(3286)

Geometry: Floor plans for a building have the four corners of a room located at the points (2,3)...A
, (11,6)....B
, (-3,18)....C
, and (8,21)...D
. Determine whether the side through the points (2,3)....A
and (11,6).....B
is perpendicular ot the side through the points (2,3).....A
and (-3,18)...........C.
SLOPE OF AB = (6-3)/(11-2)=3/9=1/3
SLOPE OF AC = (18-3)/(-3-2)=15/-5=-3
PRODUCT OF SLOPES OF AB AND AC =(1/3)(-3)=-1....
HENCE THE 2 LINES ARE PERPENDICULAR.